Solve each of the following systems by addition. If a unique solution does not exist, state whether the system is consistent or dependent.
x + 5y = 10
-2x-10y=-10

Answer by venugopalramana(3286)

Solve the following system by addition. If a unique solution does not exist, state
whether the system is inconsistent or dependent.
x + 5y = 10...................................................I
-2x – 10y = -20.................................II
-2*EQN.I GIVES
-2X-10Y=-20....WHICH IS SAME AS EQN.II...
HENCE THEY ARE DEPENDENT EQUATIONS.
SO INEFFECT WE HAVE ONLY ONE EQN.TO SOLVE FOR 2 UNKOWNS.
HENCE WE SHALL HAVE INFINTE SOLUTIONS LIKE IN THIS CASE
X=0....Y=2
X=10...Y=0...ETC..
IMPORTANT NOTE.
THEY HAVE INFINITE SOLUTIONS BUT NOT ANY AND EVERY VALUE OF X AND Y.FOR EXAMPLE IF X=0 ,THEN Y HAS TO EQUAL 2 ONLY NOT ANY OTHER VALUE.SO IT HAS INFINITE SETS OF X AND Y PAIRS FOR ITS SOLUTION